The width of a rectangular gate is 2 meters larger than its height. The diagonal brace measures square root of 6 meters. Find the width and height.
Answer by nerdybill(7384)

Draw a diagram -- label what you know. You'll see that the length, width and diagonal forms a right triangle. If so, then you can apply Pythagorean theorem:
.
Let h = height
then h+2 = width
from Pythagorean we have
h^2 + (h+2)^2 = 6
h^2 + h^2 + 4h + 4 = 6
2h^2 + 4h + 4 = 6
2h^2 + 4h - 2 = 0
divide both sides by 2:
h^2 + 2h - 1 = 0
.
Since we can't factor, you must use the quadratic equation which will yield:
h = 0.414 meters (height)
h+2 = 2.414 meters (width)